# Characterization of flip process rules with the same trajectories

### v2

@@ 3.3. Proofs of lemmas
[…]
@@
In this section we prove Theorem 1.4. We briefly describe our proof strategy here. Theorem 2.3 recasts the question of whether two rules of the same order have the same trajectories in terms of the quantities $q_{J,\mathcal{R}}$, so understanding these quantities is key. On the one hand, the presence of indicator functions $\mathbb{I}_{\left\{ab\in E(H)\right\}}$ in the general expression for $q_{J,\mathcal{R}}$ given by (3) makes it difficult to work with directly. On the other hand, we observe that the symmetries present in symmetric rules allow us to obtain cleaner expressions for the quantities $q_{J,\mathcal{R}}$; this is the focus of Lemma 4.2. Indeed, we can directly verify the ()($\Leftarrow$)( ⇐ ) direction by combining Theorem 2.3 with the simplified expressions from this key lemma. For the ()($\Rightarrow$)( ⇒ ) direction, the simplified expressions from Lemma 4.2 gives us enough leeway and control to construct ‘perturbed’ versions of symmetric non-deterministic rules of order $k\geq 3$ which have the same trajectories. To handle non-symmetric rules and complete the proof, we observe and directly verify that every such rule has a natural symmetric version with the same trajectories. This gives our characterization of unique rules.
To prepare for our proof of Theorem 1.4, we introduce some notation, terminology and useful lemmas. The proof of Theorem 1.4 is given in Section 4.1. We use the language of group actions to describe the symmetries observed. Let a group $\Gamma$ act on a set $X$. For $x\in X$ the orbit of $x$ is $\operatorname{Orb}_{\Gamma}(x)=\{\gamma\cdot x:\gamma\in\Gamma\}$ and the stabilizer of $x$ is $\operatorname{Stab}_{\Gamma}(x)=\{\gamma\in\Gamma:\gamma\cdot x=x\}$. Set $\operatorname{ORB}_{\Gamma}(X)=\{\operatorname{Orb}_{\Gamma}(x):x\in X\}$. Let $\Sigma$ be a subgroup $\Gamma$. A left coset of $\Sigma$ in $\Gamma$ is a set $\gamma\Sigma=\{\gamma\sigma:\sigma\in\Sigma\}$ with $\gamma\in\Gamma$. The following technical lemma allows us to recast sums over orbits of group actions as sums over groups.
@@
Now we describe the group actions we work with. In addition to the action of $S_{k}$ on $\mathcal{H}_{k}$ described just before Theorem 1.4, we also have the following actions of $S_{k}$ on $\binom{[k]}{2}$ and $[k]^{(2)}$. For $\sigma\in S_{k}$ and $ab\in\binom{[k]}{2}$ we set $\sigma\cdot ab=\sigma(a)\sigma(b)$. For $\sigma\in S_{k}$ and $(a,b)\in[k]^{(2)}$ we set $\sigma\cdot(a,b)=(\sigma(a),\sigma(b))$. For $\ell\in\mathbb{N}$ and $A_{1},\dots,A_{\ell}\in\{\mathcal{H}_{k},\binom{[k]}{2},[k]^{(2)}\}$ we have the following action of $S_{k}$ on $\mathbf{A}=\prod_{i\in[\ell]}A_{i}$. For $\sigma\in S_{k}$ and $(a_{i})_{i\in[\ell]}\in\mathbf{A}$ set $\sigma\cdot(a_{i})_{i\in[\ell]}=(\sigma\cdot a_{i})_{i\in[\ell]}$. Recall that $\mathcal{G}_{k}=\mathcal{H}_{k}\times[k]^{(2)}$. Observe that $\mathcal{J}_{k}$ is simply the set $\operatorname{ORB}_{S_{k}}(\mathcal{G}_{k})$ of orbits for the action of $S_{k}$ on $\mathcal{G}_{k}$. Note that every group action of $S_{k}$ on a set $X$ naturally induces a group action of each subgroup $\Gamma$ of $S_{k}$ on the same set $X$.
@@
Proof of Theorem 1.4 ()($\Leftarrow$)( ⇐ ).
@@
Proof of Theorem 1.4 ()($\Rightarrow$)( ⇒ ).
We prove the contrapositive statement. We first consider when $\mathcal{R}$ is not symmetric. Set $\mathcal{A}=\mathcal{H}_{k}\times\mathcal{H}_{k}$ and $\mathcal{B}=\operatorname{ORB}_{S_{k}}(\mathcal{A})$. The symmetric version of $\mathcal{R}$ is the matrix ×$\mathcal{R}^{\times}=(\mathcal{R}^{\times}_{F,H})_{F,H\in\mathcal{H}_{k}}$ where for all $(F,H)\in B\in\mathcal{B}$ we have $\mathcal{R}^{\times}_{F,H}=\frac{1}{|B|}\sum_{(F^{\prime},H^{\prime})\in B}\mathcal{R}_{F^{\prime},H^{\prime}}$. The following claim shows that $\mathcal{R}^{\times}$ is a symmetric rule of the same order with the same trajectories as $\mathcal{R}$; since $\mathcal{R}$ is not symmetric, this shows that $\mathcal{R}$ is not unique.
@@
The only rule of order $1$ is symmetric and deterministic, so it remains to consider when $\mathcal{R}$ is a non-deterministic symmetric rule of order $k\geq 3$. It is straightforward to check that a symmetric rule $\mathcal{R}$ is deterministic if and only if for all $F\in\mathcal{H}_{k}$ and $I\in B_{F}$ there exists $q_{F,I}\in\{0,|I|\}$ such that for all $H\in\mathcal{H}_{k}$ such that $p^{F,H}_{I}\neq q_{F,I}$ we have $\mathcal{R}_{F,H}=0$. Hence, the following (not necessarily mutually exclusive) cases cover all non-deterministic symmetric rules of order $k\geq 3$.
@@ 5.1. Characterization by flip process distributions
    For all $B\in\operatorname{ORB}_{S_{k}}(\mathcal{H}_{k}\times\mathcal{H}_{k})$ we have $\sum_{(F^{\prime},H^{\prime})\in B}\mathcal{R}_{F^{\prime},H^{\prime}}=\sum_{(F^{\prime},H^{\prime})\in B}\mathcal{R}^{\prime}_{F^{\prime},H^{\prime}}$

### v3

@@ 3.3. Proofs of lemmas
We introduce some further definitions. Let $\mathbf{F}$ and $\mathbf{H}$ be rooted graphs. A function $\phi\colon V(\mathbf{F})\to V(\mathbf{H})$ is relation-preserving if it satisfies (F1), root-respecting if it satisfies (F2), and root-order-preserving if it satisfies (F3). In particular, an induced copy function between rooted graphs is relation-preserving, root-respecting and root-order-preserving.
@@
For $\mathbf{x}\in\Omega^{V(\mathbf{F}(\mathbf{m}))}$ define $f_{\mathbf{x}}\colon V(\mathbf{F}(\mathbf{m}))\to V(\mathbf{G}^{\times})$ as follows. For $i\in V(\mathbf{F}(\mathbf{m}))$ set $f_{\mathbf{x}}(i)=w$ where $x_{i}\in\Omega_{w}$. Let $(p,q)=R(\mathbf{F}(\mathbf{m}))$. Set $\mathcal{P}=\{\mathbf{x}\in\Omega^{V(\mathbf{F}(\mathbf{m}))}:f_{\mathbf{x}}\textrm{ is relation-preserving}\}$ and $\mathcal{Q}=\{\mathbf{x}\in\mathcal{P}:x_{p}=x,x_{q}=y\}$. For $\mathbf{x}\in\mathcal{Q}$ define functions $g_{\mathbf{x}}\colon V(\mathbf{F}(\mathbf{m}))\to V(\mathbf{H})$ and $h_{\mathbf{x}}\colon V(\mathbf{F})\to V(\mathbf{H})$ as follows. For $i\in V(\mathbf{F}(\mathbf{m}))$ set
@@
In this section we prove Theorem 1.4. We briefly describe our proof strategy here. Theorem 2.3 recasts the question of whether two rules of the same order have the same trajectories in terms of the quantities $q_{J,\mathcal{R}}$, so understanding these quantities is key. On the one hand, the presence of indicator functions $\mathbb{I}_{\left\{ab\in E(H)\right\}}$ in the general expression for $q_{J,\mathcal{R}}$ given by (3) makes it difficult to work with directly. On the other hand, we observe that the symmetries present in symmetric rules allow us to obtain cleaner expressions for the quantities $q_{J,\mathcal{R}}$; this is the focus of Lemma 4.2. Indeed, we can directly verify the $(\Leftarrow)$ direction of Theorem 1.4 by combining Theorem 2.3 with the simplified expressions from this key lemma. For the $(\Rightarrow)$ direction, the simplified expressions from Lemma 4.2 give us enough leeway and control to construct ‘perturbed’ versions of symmetric non-deterministic rules of order $k\geq 3$ which have the same trajectories. To handle non-symmetric rules and complete the proof, we observe and directly verify that every such rule has a natural symmetric version with the same trajectories. This gives our characterization of unique rules.
To prepare for our proof of Theorem 1.4, we introduce some notation, terminology and useful lemmas. The proof of Theorem 1.4 is given in Section 4.1. We use the language of group actions to describe the symmetries observed. Let a group $\Gamma$ act on a set $X$. For $x\in X$ the orbit of $x$ is $\operatorname{Orb}_{\Gamma}(x)=\{\gamma\cdot x:\gamma\in\Gamma\}$ and the stabilizer of $x$ is $\operatorname{Stab}_{\Gamma}(x)=\{\gamma\in\Gamma:\gamma\cdot x=x\}$. Set $\operatorname{ORB}_{\Gamma}(X)=\{\operatorname{Orb}_{\Gamma}(x):x\in X\}$. Let $\Sigma$ be a subgroup of $\Gamma$. A left coset of $\Sigma$ in $\Gamma$ is a set $\gamma\Sigma=\{\gamma\sigma:\sigma\in\Sigma\}$ with $\gamma\in\Gamma$. The following technical lemma allows us to recast sums over orbits of group actions as sums over groups.
@@
Now we describe the group actions we work with. In addition to the action of $S_{k}$ on $\mathcal{H}_{k}$ described just before Theorem 1.4, we also have the following actions of $S_{k}$ on $\binom{[k]}{2}$ and $k]^{(2)}$. For $\sigma\in S_{k}$ and $ab\in\binom{[k]}{2}$ we set $\sigma\cdot ab=\sigma(a)\sigma(b)$. For $\sigma\in S_{k}$ and $(a,b)\in[k]^{(2)}$ we set $\sigma\cdot(a,b)=(\sigma(a),\sigma(b))$. For $\ell\in\mathbb{N}$ and $A_{1},\dots,A_{\ell}\in\{\mathcal{H}_{k},\binom{[k]}{2},[k]^{(2)}\}$ we have the following action of $S_{k}$ on $\mathbf{A}=\prod_{i\in[\ell]}A_{i}$. For $\sigma\in S_{k}$ and $(a_{i})_{i\in[\ell]}\in\mathbf{A}$ set $\sigma\cdot(a_{i})_{i\in[\ell]}=(\sigma\cdot a_{i})_{i\in[\ell]}$. Recall that $\mathcal{G}_{k}=\mathcal{H}_{k}\times[k]^{(2)}$. Observe that $\mathcal{J}_{k}$ is simply the set $\operatorname{ORB}_{S_{k}}(\mathcal{G}_{k})$ of orbits for the action of $S_{k}$ on $\mathcal{G}_{k}$. Note that every group action of $S_{k}$ on a set $X$ naturally induces a group action of each subgroup $\Gamma$ of $S_{k}$ on the same set $X$.
@@
Proof of Theorem 1.4 $(\Leftarrow)$.
@@
Proof of Theorem 1.4 $(\Rightarrow)$.
We prove the contrapositive statement. We first consider when $\mathcal{R}$ is not symmetric. Set $\mathcal{A}=\mathcal{H}_{k}\times\mathcal{H}_{k}$ and $\mathcal{B}=\operatorname{ORB}_{S_{k}}(\mathcal{A})$. The symmetric version of $\mathcal{R}$ is the matrix $\mathcal{R}^{\times}=(\mathcal{R}^{\times}_{F,H})_{F,H\in\mathcal{H}_{k}}$ where for all $(F,H)\in B\in\mathcal{B}$ we have $\mathcal{R}^{\times}_{F,H}=\frac{1}{|B|}\sum_{(F^{\prime},H^{\prime})\in B}\mathcal{R}_{F^{\prime},H^{\prime}}$. The following claim shows that $\mathcal{R}^{\times}$ is a symmetric rule of the same order with the same trajectories as $\mathcal{R}$; since $\mathcal{R}$ is not symmetric, this shows that $\mathcal{R}$ is not unique.
@@
The only rule of order $1$ is symmetric and deterministic, so it remains to consider when $\mathcal{R}$ is a non-deterministic symmetric rule of order $k\geq 3$. First we prove the following claim.
Claim 4.4.
A symmetric rule $\mathcal{R}$ is deterministic if and only if for all $F\in\mathcal{H}_{k}$ and $I\in B_{F}$ there exists $q_{F,I}\in\{0,|I|\}$ such that for all $H\in\mathcal{H}_{k}$ such that $p^{F,H}_{I}\neq q_{F,I}$ we have $\mathcal{R}_{F,H}=0$.
Proof.
We start with the $(\Rightarrow)$ direction. Since $\mathcal{R}$ is deterministic, we have a unique $G_{F}\in\mathcal{H}_{k}$ such that $\mathcal{R}_{F,H}=\mathbb{I}_{\left\{H=G_{F}\right\}}$. Take $F\in\mathcal{H}_{k}$ and $I\in B_{F}$. We set $q_{F,I}:=|E(G_{F})\cap I|$ and claim that $q_{F,I}\in\{0,|I|\}$. Indeed, suppose that $q_{F,I}>0$. Then we may pick $e\in E(G_{F})\cap I$. Now take $\sigma\in S_{F}$. Since $\mathcal{R}$ is symmetric, we have $1=\mathcal{R}_{F,G_{F}}=\mathcal{R}_{\sigma\cdot F,\sigma\cdot G_{F}}=\mathcal{R}_{F,\sigma\cdot G_{F}}$, which implies $G_{F}=\sigma\cdot G_{F}$. In particular, we have $\sigma\cdot e\in E(G_{F})$. Since $I=\operatorname{Orb}_{S_{F}}(e)$, we have $E(G_{F})\cap I=I$ as required. Now take $H\in\mathcal{H}_{k}$ such that $p^{F,H}_{I}=|E(H)\cap I|\neq|E(G_{F})\cap I|=q_{F,I}$. In particular, we have $H\neq G_{F}$. Since $\mathcal{R}$ is deterministic, we have $\mathcal{R}_{F,H}=0$.
Now we consider the $(\Leftarrow)$ direction. Fix $F\in\mathcal{H}_{k}$ and for all $I\in B_{F}$ take $q_{F,I}\in\{0,|I|\}$ whose existence is guaranteed by assumption. Let $G_{F}\in\mathcal{H}_{k}$ be the labelled graph on $[k]$ such that $E(G_{F})=\bigcup_{I:q_{F,I}=|I|}I$. Now $G_{F}$ is the unique labelled graph $H\in\mathcal{H}_{k}$ such that $p^{F,H}_{I}=|E(H)\cap I|=q_{F,I}$ for all $I\in B_{F}$, so for all $H\in\mathcal{H}_{k}$ we have $\mathcal{R}_{F,H}=\mathbb{I}_{\left\{H=G_{F}\right\}}$. It follows that $\mathcal{R}$ is deterministic. ∎
It follows from Claim 4.4 that the following (not necessarily mutually exclusive) cases cover all non-deterministic symmetric rules of order $k\geq 3$.
@@ 5.1. Characterization by flip process distributions
    For all $B\in\operatorname{ORB}_{S_{k}}(\mathcal{H}_{k}\times\mathcal{H}_{k})$ we have $\sum_{(F^{\prime},H^{\prime})\in B}\mathcal{R}_{F^{\prime},H^{\prime}}=\sum_{(F^{\prime},H^{\prime})\in B}\mathcal{R}^{\prime}_{F^{\prime},H^{\prime}}$ .